In 5 minutes, a current of 5 amps is established in a wire. How many coulombs and how many electrons pass through each cross section of this wire?


1
Expert's answer
2021-04-26T16:48:33-0400

The current is

"\displaystyle I = \frac{q}{t} \; \Longrightarrow \; q = It"

"q = 5 A \cdot5 \cdot 60 s = 1500\,C"

One electron has charge "1.6 \cdot 10^{-19} \, C". Then

"\displaystyle N = \frac{q}{1.6 \cdot 10^{-19}} = \frac{1500}{1.6 \cdot 10^{-19}} = 937.5 \cdot 10^{19} = 9.375 \cdot 10^{21}"

Answer: "1500\; C, \, 9.375 \cdot 10^{21}"